# Numerical Twin with Two Dimensional Ornstein–Uhlenbeck Processes of Transient Oscillations in EEG signal

### v1

@@ V Non-stationary OU-process
where $m_{i}$ is the mean of the signal on segment i. When $\sigma_{T}(i)>T_{std}$ and $mask(i)>T_{m}$, we label such segment $i$ as non-stationary. We found several stationary segment, where we can estimated the parameters of the OU-process (fig. 5B). To conclude, the present decomposition on a OU-process can be computed for non-stationnary processes, under the condition that there are stationary epochs that are long enough.
Figure 5: Non-stationary Ornstein–Uhlenbeck model. (A) Dynamics parameters $\lambda$ (decay rate), $\omega$ (rotation frequency), and $\sigma$ (noise amplitude) of the two-dimensional Ornstein–Uhlenbeck (OU) model over time. The parameters are estimated sequentially from EEG recordings filtered in the $\alpha$-band. (B) Piecewise approximation of the parameter trajectories using constant segments, separated by detected transition points (black vertical lines). This segmentation captures long-range trends and identifies parameter regimes with relatively stable dynamical behavior. (C) Time-frequency representation (spectrogram) of the EEG signal, showing spindle activity evolving over the course of the recording. (D) Raw signal (gray) and smoothed EEG (blue) used for parameter extraction. Together, these panels illustrate the feasibility of modeling EEG dynamics with a time-varying stochastic focus model, providing a numerical twin of the patient’s brain state during anesthesia or sleep.
We can now generalize the non-stationary representation to the case of multiple dominant frequency bands so that we decompose the EEG signal as a sum o
@@
where $\text{ou}_{k}$ are solutions of eq. 1 with parameters $\bar{\lambda_{k}},\bar{\omega}_{k},\bar{\sigma}_{k}$ be fitted to data and $\text{res}(t)$ is a residual component. This is equivalent to the power spectral decomposition identity in a sliding window [t,t+T]
@@
where the three parameters $(\lambda(t),\omega(t),\sigma(t))$ are piece-wise constant and are estimated for each band computed on sliding windows of few minutes to obtain a local steady-state estimation, when the statistics do not vary much in time (Fig. 5). The residual term $\text{Res}(f,t)$ that accounts for the general trend of power decay in the EEG
@@ VI Discussion
Short-time Fourier and wavelet transforms [Mey90I] offer excellent spectral localization but do not by themselves produce a compact, mechanistic model for sporadic, irregular bursts with slow envelope modulation. Empirical Mode Decomposition (EMD) and related envelope-based approaches [flandrin2004empirical] capture instantaneous amplitudes and phases but remain descriptive. Our framework bridges these viewpoints: it uses envelope information to segment events yet constrains the dynamics with a linear stochastic focus model that yields closed-form autocorrelation and PSD expressions, making estimation principled and interpretable. Methods that separate $1/f$-aperiodic backgrounds from peaks (e.g., IRASA and parametric spectral fits [wen_separating_2016, donoghue2020parameterizing]) are complementary; here, such backgrounds are treated as residual components outside the OU bands, while peaks are fitted by OU Lorentzians with physically meaningful parameters.
@@ Acknowledgments
D.H. research is supported by ANR grants AstroXcite and AnalysisSpectralEEG and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 882673).
@@
The solution $x(t)$ or $y(t)$ of Eq. (1) is computed using a fourth-order Runge-Kutta scheme (Fig. 1C). For the range of considered parameters, spindles are generated.
@@
On Weak Coupling and Cross-Correlation
Two uncoupled bands modeled by independent OU foci are uncorrelated. With small linear coupling $\epsilon$, cross-correlation terms are $O(\epsilon)$ and decay as $e^{-\lambda|\tau|}$, with amplitude diminishing rapidly for widely separated $\omega$’s. Consider the system of coupled Ornstein–Uhlenbeck processes:

### v2

@@ V Non-stationary OU-process
where $m_{i}$ is the mean of the signal on segment i. When $\sigma_{T}(i)>T_{std}$ and $mask(i)>T_{m}$, we label such segment $i$ as non-stationary. We found several stationary segment, where we estimated the parameters of the OU-process (fig. 5B). To conclude, the present decomposition on a OU-process can be computed for non-stationnary processes, under the condition that there are stationary epochs that are long enough.
Figure 5: Non-stationary Ornstein–Uhlenbeck model. (A) Dynamics parameters $\lambda$ (decay rate), $\omega$ (rotation frequency), and $\sigma$ (noise amplitude) of the two-dimensional Ornstein–Uhlenbeck (OU) model over time. The parameters are estimated sequentially from EEG recordings filtered in the $\alpha$-band. (B) Piecewise approximation of the parameter trajectories using constant segments, separated by detected transition points (black vertical lines). This segmentation captures long-range trends and identifies parameter regimes with relatively stable dynamical behavior. (C) Time-frequency representation (spectrogram) of the EEG signal, showing spindle activity evolving over the course of the recording. (D) Raw signal (gray) and smoothed EEG (blue) used for parameter extraction. Together, these panels illustrate the feasibility of modeling EEG dynamics with a time-varying stochastic focus model (or equivalently linear-Gaussian continuous-time oscillator), providing a numerical twin of the patient’s brain state during anesthesia or sleep.
We can now generalize the non-stationary representation to the case of multiple dominant frequency bands so that we decompose the EEG signal as a sum
@@
where $\text{ou}_{k}$ are solutions of Eq. 1 with parameters $\bar{\lambda_{k}},\bar{\omega}_{k},\bar{\sigma}_{k}$ fitted to the data and $\text{res}(t)$ is a residual component. This is equivalent to the power spectral decomposition identity in a sliding window [t,t+T]
@@
where the three parameters $(\lambda(t),\omega(t),\sigma(t))$ are piecewise constant and are estimated for each band computed on sliding windows of few minutes to obtain a local steady-state estimation, when the statistics do not vary much in time (Fig. 5). The residual term $\text{Res}(f,t)$ accounts for the general trend of power decay in the EEG
@@ VI Discussion
Short-time Fourier and wavelet transforms [19] offer excellent spectral localization but do not by themselves produce a compact, mechanistic model for sporadic, irregular bursts with slow envelope modulation. Empirical Mode Decomposition (EMD) and related envelope-based approaches [21] capture instantaneous amplitudes and phases but remain descriptive. Our framework bridges these viewpoints: it uses envelope information to segment events yet constrains the dynamics with a linear stochastic focus model that yields closed-form autocorrelation and PSD expressions, making estimation principled and interpretable. Methods that separate $1/f$-aperiodic backgrounds from peaks (e.g., IRASA and parametric spectral fits [32, 33]) are complementary; here, such backgrounds are treated as residual components outside the OU bands, while peaks are fitted by OU Lorentzians with physically meaningful parameters.
@@ Acknowledgments
We confirm that EEG were recorded from informed consent patients. D.H. research is supported by ANR grants AstroXcite and AnalysisSpectralEEG and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 882673). We use Chatgpt5.2 for correcting some sentences and verifying some computations.
@@
The solution $x(t)$ or $y(t)$ of Eq. (1) is computed using a Euler’s stochastic scheme (Fig. 1C). For the range of considered parameters, spindles are generated.
@@
Weak Coupling and Cross-Correlation
Two uncoupled bands modeled by independent OU foci are uncorrelated. With small linear coupling $\epsilon$, cross-correlation terms are $O(\epsilon)$ and decay as $e^{-\lambda|\tau|}$, with amplitude decreasing rapidly for widely separated $\omega$’s. Consider the system of coupled Ornstein–Uhlenbeck processes: